OpenStudy (anonymous):

what are the molarity and mole fraction of solute in a 35.5 percent by mass aqueous solution of formic acid (HCOOH}?

OpenStudy (ria23):

35.5 g of HCOOH in 64.5 g of water Moles HCOOH = 35.5 g / 30.02 g/mol = 1.18 moles water = 64.5 g / 18.02 g/mol = 3.58 Mole fraction = 1.18 / 1.18 + 3.58 = 0.248 m = 1.18 / 0.0645 Kg = 18.3 There yhu go. C:

OpenStudy (anonymous):

@Ria23 thank you so much for your help

OpenStudy (ria23):

Yhur welcome. C: